OpenStudy (anonymous):

Suppose U = {1,2,3,4,5,6,7,8,9,10} is the universal set and Q = {3,6,9} What is Q?

OpenStudy (anonymous):

What is Q? Well they gave it. Do they want Q or Q compliment?

OpenStudy (anonymous):

Q compliment

OpenStudy (anonymous):

Just take all the numbers in U set that aren't in Q set.

OpenStudy (anonymous):

You can do it.

OpenStudy (anonymous):

Still confused?

OpenStudy (anonymous):

Very much so

OpenStudy (anonymous):

Oh! I understand!

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

So, wait, would it be 1,4,5,7,8,9,10 etc?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Not quite, let me show.

OpenStudy (anonymous):

\[ \overline{Q}= U-Q= \{1,2,\cancel3,4,5,\cancel6,7,8,\cancel9,10\} - \{\cancel3,\cancel6,\cancel9\} = \{1,2,4,5,10\} \]

OpenStudy (anonymous):

Ahh, okay. Thank you @wio I have one more question for you

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Suppose U = {1,2,3,4,5,6,7,8} and A = {1,3,5,7} and B = {4,5,6} A subset B True or False? B subset U True or False? @wio

OpenStudy (anonymous):

Well, if A is a subset of B, then \(A-B=\emptyset=\{\}\)

OpenStudy (anonymous):

So you can use set different to tell.

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Do you know what a subset is?

OpenStudy (anonymous):

To be honest, no.

OpenStudy (anonymous):

That is the problem, you need to understand it. If A is a subset of B, for example, that means that every element in A is also in B

OpenStudy (anonymous):

Oh, okay, so then A subset B is false

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Yeah, it is false.

OpenStudy (anonymous):

But, B subset U isn't because {4,5,6} are all in {1,2,3,4,5,6,7} @wio

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

B is a subset of U

OpenStudy (anonymous):

Thank you!